# Investigating the Event-Shape Methods in Search for the Chiral Magnetic Effect in Relativistic Heavy Ion Collisions

### v3

@@ Abstract
The Chiral Magnetic Effect (CME) is a phenomenon in which electric charge is separated by a strong magnetic field from local domains of chirality imbalance and parity violation in quantum chromodynamics (QCD). The CME-sensitive observable, charge-dependent three-point azimuthal correlator $\Delta\gamma$, is contaminated by a major physics background proportional to the particle’s elliptic flow anisotropy $v_{2}$. Event-shape engineering (ESE) binning events in dynamical fluctuations of $v_{2}$ and event-shape selection (ESS) binning events in statistical fluctuations of $v_{2}$ are two methods to search for the CME by projecting $\Delta\gamma$ to the measured anisotropy $v_{2}=0$ intercept. We conduct a systematic study of these two methods using physics models as well as toy model simulations. It is observed that the ESE method fulfills the general premise of measuring the CME but is statistically hungry. It is found that the intercept from the ESS method depends on the details of the event content, such as the mixtures of background-contributing sources, because of statistical fluctuations of intertwining variables used in the method, and is thus not practically useful or clean to measure the CME.
@@ I Introduction
It has been predicted by quantum chromodynamics (QCD) that vacuum fluctuations can result in gluon fields of nonzero topological charges in local metastable domains. Interactions of quarks with such gluon fields can cause chirality imbalance, breaking the parity and charge-parity symmetries in those domains. Such chirality imbalance, under a strong magnetic field, would result in charge separation, a phenomenon called the chrial magnetic effect (CME) [1, 2, 3, 4].
Ultra-strong magnetic fields are presumably produced in non-head-on relativistic heavy ion collisions [5, 6]. The magnetic field is on average perpendicular to the reaction plane (RP, the plane span by the beam and the impact parameter direction of the collision). The CME-induced charge separation is along the direction of the magnetic field, and may be conveniently quantified by the $a_{1}$ variable in Fourier series of particle azimuthal distributions [7, 8],
$dN_{\pm}/d\tilde{\phi}\propto 1\pm 2a_{1}\sin\tilde{\phi}+2v_{2}\cos 2\tilde{\phi}+\cdots\,,$ (1)
where $\tilde{\phi}$ is the azimuthal angle of the particle momentum vector with respect to the RP and the subscript ‘$\pm$’ indicates particle charge sign. The elliptic anisotropy $v_{2}$ harmonic is the leading modulation in particle distributions produced in relativistic heavy ion collisions [9]. Because of a vanishing mean $a_{1}$ due to random fluctuations, a commonly used observable is the three-point correlator [8],
@@ II The Event-Shape Techniques
In ESE, the $q_{2}^{2}$ is typically measured at forward and backward rapidities [29, 30]. The event classes selected by $q_{2}$ (or equivalently $q_{2}^{2}$ or $\hat{q}_{2}^{2}$) within a given narrow centrality range will have a different average $\left\langle v_{2}\right\rangle$ in the midrapidity region, or any other kinematic regions displaced from that for $q_{2}^{2}$ computation. In other words, because $q_{2}$ is correlated with $\left\langle v_{2}\right\rangle$, ESE is dividing events in groups of varying $\left\langle v_{2}\right\rangle$, which dynamically fluctuates from event to event. These dynamical fluctuations are mainly caused by initial geometry fluctuations within the given narrow centrality bin [41, 42].
@@
Since the ESE method selects events according to dynamical fluctuations of the $\left\langle v_{2}\right\rangle$ of the POIs, decoupled from the $q_{2}^{2}$ computation, it is guaranteed that the average resonance $\left\langle v_{\rm 2,{\rm res}}\right\rangle$, or generally the elliptic anisotropies of all elements in the event including all CME background sources, are proportional to $\left\langle v_{2}\right\rangle$. Note that the final-state $\left\langle v_{2}\right\rangle$ contains contributions not only from primordial particles, but also from all decay products of resonances and clusters. Since the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and $\left\langle v_{2}\right\rangle$ in those $q_{2}$-selected event classes are all connected to the initial geometry, they are all proportional to each other. This is illustrated in the cartoon in Fig. 1 (left), where the three ellipses illustrate event selections by $q_{2}$ and the hallow circles are the average $\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ in each $q_{2}$-selected event class. In the ESE analyses for CME searches [29, 30], one analyzes $\Delta\gamma$ as a function of $\left\langle v_{2}\right\rangle$ in each event class and projects the $\Delta\gamma$ measurement to zero $\left\langle v_{2}\right\rangle$. The ESE method guarantees that the elliptic flows of the background sources also vanish at zero $\left\langle v_{2}\right\rangle$. In other words, a linear projection to $\left\langle v_{2}\right\rangle=0$ would be equivalent to a projection to $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$, where the CME backgrounds are zero. Since all those events are from a narrow centrality bin where the CME signal does not vary much, the intercept of the linear projection is presumably the CME signal. Note that there may still be complications from nonflow contamination [43] depending on how $\Delta\gamma$ and $\left\langle v_{2}\right\rangle$ are measured, which is outside the scope of the present work.
@@
Since the CME background comes from the elliptic anisotropies imprinted in the POIs, it is interesting to select events with an event-shape ellipticity $v_{2}^{\rm obs}\equiv\left\langle\cos 2(\phi-\psi)\right\rangle$ calculated by POIs themselves (where the EP $\psi$ is determined in another kinematic region away from the POIs) and examine the $\Delta\gamma$ observable as a function of $v_{2}^{\rm obs}$. This method selects on the event-by-event elliptic-shape quantity (particle emissin pattern) $v_{2}^{\rm obs}$ which can fluctuate even to negative values because of large statistical fluctuations, the intercept at $v_{2}^{\rm obs}=0$, more sensitive to the CME, can be well determined from data. This event-shape (ES) analysis was performed by STAR [24] and the intercept of the linear fit is consistent with zero with the then-available statistics. However, the backgrounds are not determined by $v_{2}^{\rm obs}$ but by $\left\langle v_{\rm 2,{\rm res}}\right\rangle$. Statistically fluctuated $v_{2}^{\rm obs}=0$ does not necessarily guarantee $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$. This is illustrated by the cartoon in Fig. 1 (center). Although $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ is proportional to $\left\langle v_{2}\right\rangle$ with an unbiased event selection, the strongly (maximally) biased selection of events by $v_{2}^{\rm obs}$ will obscure the relationship of the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and $v_{2}^{\rm obs}$ in each bin of $v_{2}^{\rm obs}$ of those events, as indicated by the hollow box in Fig. 1 (center). Because of the overall positive means $\left\langle v_{2}^{\rm obs}\right\rangle=\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ over the entire event sample, the linear projection will give a positive intercept of $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ at $v_{2}^{\rm obs}=0$. In other words, with finite positive average $\left\langle v_{2}\right\rangle$ in heavy ion collisions, $\left\langle v_{\rm 2,{\rm res}}\right\rangle>0$ at $v_{2}^{\rm obs}=0$ with events binned in $v_{2}^{\rm obs}$. This has been verified by model study [44]. Thus, positive residual background remains in the linear fit intercept in the analysis of $\Delta\gamma$ as a function of $v_{2}^{\rm obs}$ [24], although the background is significantly reduced compared to that in the inclusive $\Delta\gamma$ measurement. Ideally, in the ES analysis, one would want to select events with the $v_{2}^{\rm obs}$ values of all resonances (background sources) to be zero, however, such selection is experimentally insurmountable.
@@
It has been recently proposed by Xu et al. [36] to select events according to the $\hat{q}_{2,\textsc{poi}}^{2}$ of the POIs, and study the $\Delta\gamma$ as a function of $v_{2}$ of the POIs. This is similar in spirit to the ES method above where the events are selected by $v_{2}^{\rm obs}$ and the $\Delta\gamma$ is examined as a function of the same $v_{2}^{\rm obs}$ variable. Here, the selection variable $\hat{q}_{2,\textsc{poi}}^{2}$ and the $v_{2}$ variable are different, although related as they are computed by the same POIs. The $\hat{q}_{2}^{2}$ variable is an event-by-event quantity connected to two-particle cumulant anisotropy (cf Eq. 5), whereas $v_{2}$ can be measured with respect to $\psi$ from another kinematic region like in the ES method. (Of course $v_{2}$ can also be computed by the two-particle cumulant method, in which case it is more directly connected to $\hat{q}_{2}^{2}$.) The variations in the event-by-event $\hat{q}_{2,\textsc{poi}}^{2}$ quantity, within a given narrow centrality bin, are mainly from statistical fluctuations. Since the $v_{2}$ in those $\hat{q}_{2,\textsc{poi}}^{2}$-selected events are related to $\hat{q}_{2,\textsc{poi}}^{2}$ itself, the variations in $v_{2}$ is also primarily of statistical nature. (Note, we simply use $v_{2}$ here, instead of the $\left\langle v_{2}\right\rangle$ for ESE, to indicate that the $v_{2}$ in ESS is of statistical nature although it is an “average” $v_{2}$ in events selected by $\hat{q}_{2,\textsc{poi}}^{2}$.) This technique selects on the statistical fluctuations in event shape, not on the dynamical fluctuations in $\left\langle v_{2}\right\rangle$ like in the ESE method, and it is referred to as the Event-Shape-Selection (ESS) method [36]. The most important difference in the ESS method from the ESE method is that the $\hat{q}_{2}^{2}$ is computed from the POIs in the former but in the later the phase-space where to compute $\hat{q}_{2}^{2}$ is displaced from that of the POIs’. In both approaches, one groups events (within a given narrow centrality bin) according to $\hat{q}_{2}^{2}$ and studies the $\Delta\gamma$ variable as a function of the $v_{2}$ of the POIs.
@@ III Model setup and analysis details
Figure 2: $m_{\rm inv}$ distributions. Shown are the per-event $m_{\rm inv}$ distribution of OS pairs subtracted by that of SS pairs from avfd with axial current densities $n_{5}/s=0,0.1$, and $0.2$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I simulation including primordial pions, and $K_{S}$ and $\rho^{0}$-resonance decay pions (lower center), and Toy Model II simulation including in addition decay pions from a mass continuum (lower right). Particles used to form the $m_{\rm inv}$ distributions are taken from all charged hadrons (defined to be $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ in our study and treated as pions with pion mass) and are from the kinematic range of $|\eta|<1$ and $20.2<p_{\rm T}<$ 2 GeV/$c$. In the toy models all final-state particles are charged pions.
@@ III.5 Toy Model I
Two versions of the toy model are examined. In Toy Model I, we include primordial pions and two resonances, $K_{S}$ and $\rho^{0}$. The resonance mass is generated according to the Breit-Wignar distribution of proper mass and width [62]. We generate particles uniformly within the rapidity range of $-1.5<y<1.5$, taking event-by-event particle multiplicities according to Poisson statistics about the means of three times the measured average pseudorapidity densities [63, 45, 64]. The $p_{\rm T}$ spectra are parameterized based on the measured data [65, 66], namely,
@@ III.6 Toy Model II
As shown in Fig. 2, all physics models have a mass continuum in the $m_{\rm inv}$ distribution on which the resonance peaks reside. To mimic this, we add a mass continuum in our second version of the toy model. The mass distribution of the continuum is generated according to a probability $\propto e^{-m_{\rm inv}c^{2}/{\rm GeV}}$ between the two-pion mass threshold (0.28 GeV/$c^{2}$) and 5 GeV/$c^{2}$. Given a generated mass value, a Breit-Wignar distribution (mean at the mass value and width equal to the $\rho^{0}$ width [62]) is sampled to obtain the mass of the continuum “resonance.” All particles in the generated mass continuum are decayed into $\pi^{+}\pi^{-}$ isotropically in the parent rest frame. The decay pions are properly boosted to the lab frame.

### v4

@@ Abstract
The Chiral Magnetic Effect (CME) is a phenomenon in which electric charge is separated by a strong magnetic field from local domains of chirality imbalance and parity violation in quantum chromodynamics (QCD). The CME-sensitive observable, charge-dependent three-point azimuthal correlator $\Delta\gamma$, is contaminated by a major physics background proportional to the particle’s elliptic flow anisotropy $v_{2}$. Event-shape engineering (ESE) binning events in dynamical fluctuations of $v_{2}$ and event-shape selection (ESS) binning events in statistical fluctuations of $v_{2}$ are two methods to search for the CME by projecting $\Delta\gamma$ to the measured anisotropy $v_{2}=0$ intercept. We conduct a systematic study of these two methods using physics models as well as toy model simulations. It is observed that the ESE method fulfills the general premise of measuring the CME but is statistically hungry. It is found that the intercept from the ESS method depends on the details of the event content, such as the mixtures of background-contributing sources, because of statistical fluctuations of intertwining variables used in the method, and is thus not practically useful to measure the CME.
@@ I Introduction
It has been predicted by quantum chromodynamics (QCD) that vacuum fluctuations can result in gluon fields of nonzero topological charges in local metastable domains. Interactions of quarks with such gluon fields can cause chirality imbalance, breaking the parity and charge-parity symmetries in those domains. Such chirality imbalance, under a strong magnetic field, would result in charge separation, a phenomenon called the chiral magnetic effect (CME) [1, 2, 3, 4].
Ultra-strong magnetic fields are presumably produced in non-central relativistic heavy ion collisions [5, 6]. The magnetic field is on average perpendicular to the reaction plane (RP, the plane span by the beam and the impact parameter direction of the collision). The CME-induced charge separation is along the direction of the magnetic field, and may be conveniently quantified by the $a_{1}$ variable in Fourier series of particle azimuthal distributions [7, 8],
$dN_{\pm}/d\tilde{\phi}\propto 1\pm 2a_{1}\sin\tilde{\phi}+2v_{2}\cos 2\tilde{\phi}$$\cdots\,$, (1)
where $\tilde{\phi}$ is the azimuthal angle of the particle momentum vector with respect to the RP and the subscript $\pm$’ indicates particle charge sign. The elliptic anisotropy $v_{2}$ harmonic is the leading modulation in particle distributions produced in relativistic heavy ion collisions [9]. Because of a vanishing mean $a_{1}$ due to random fluctuations, a commonly used observable is the three-point correlator [8],
@@ II The Event-Shape Techniques
In ESE, the $q_{2}^{2}$ is typically measured at forward and backward rapidities [29, 30], within a given narrow centrality range, and events are selected according to the measured value of $q_{2}^{2}$ to study $\Delta\gamma$ as a function of $v_{2}$ typically in the midrapidity region. The value of $q_{2}^{2}$ varies from event to event, and the variations are dominated by statistical fluctuations. These statistical fluctuations in $q_{2}^{2}$ at forward and backward rapidities are uncorrelated to those in $v_{2}$ (or any other quantities) at midrapidity, so the statistical fluctuation effect in $v_{2}$ is averaged to zero in those $q_{2}^{2}$-selected events. There are, however, also dynamical fluctuations affecting the value of $q_{2}^{2}$. The main source of dynamical fluctuations is the initial geometry, which can fluctuate event-by-event within a given (narrow) centrality bin [41, 42], or even at a centrality precisely determined by a particular centrality measure. Even if the colliding geometry was precisely fixed, there can still be fluctuations in the subsequent “hydrodynamic” evolution and interactions of the collision system. Because the sources of these dynamical fluctuations are common to the entire collision event, they affect $q_{2}^{2}$ and $v_{2}$ (and any other quantities) of the same event. Thus, the event classes selected by $q_{2}$ (or equivalently $q_{2}^{2}$ or $\hat{q}_{2}^{2}$) within a given narrow centrality range will have a different average $\left\langle v_{2}\right\rangle$ in the midrapidity region, or any other kinematic regions displaced from that for $q_{2}^{2}$ computation. In other words, because $q_{2}$ is dynamically correlated with $\left\langle v_{2}\right\rangle$, ESE is dividing events in groups of varying $\left\langle v_{2}\right\rangle$, which dynamically fluctuates from event to event. We note that since $\hat{q}_{2}^{2}$ and $\left\langle v_{2}\right\rangle$ are usually computed in separate $\eta$ regions, longitudinal decorrelation of flow [43, 44] can reduce the power of $\hat{q}_{2}^{2}$ in selecting dynamical fluctuations in $\left\langle v_{2}\right\rangle$. However, this effect is small because the flow decorrelation has been measured to be at most a few percent [45, 46].
@@
Since the ESE method selects events according to dynamical fluctuations of the $\left\langle v_{2}\right\rangle$ of the POIs, decoupled from the $q_{2}^{2}$ computation, it is guaranteed that the average resonance $\left\langle v_{\rm 2,{\rm res}}\right\rangle$, or generally the elliptic anisotropies of all elements in the event including all CME background sources, are proportional to $\left\langle v_{2}\right\rangle$. Note that the final-state $\left\langle v_{2}\right\rangle$ contains contributions not only from primordial particles, but also from all decay products of resonances and clusters. Since the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and $\left\langle v_{2}\right\rangle$ in those $q_{2}$-selected event classes are all connected to the initial geometry, they are all proportional to each other. This is illustrated in the cartoon in Fig. 1 (left), where the three ellipses illustrate event selections by $q_{2}$ and the hallow circles are the average $\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ in each $q_{2}$-selected event class. In the ESE analyses for CME searches [29, 30], one analyzes $\Delta\gamma$ as a function of $\left\langle v_{2}\right\rangle$ in each event class and projects the $\Delta\gamma$ measurement to zero $\left\langle v_{2}\right\rangle$. The ESE method guarantees that the elliptic flows of the background sources also vanish at zero $\left\langle v_{2}\right\rangle$. In other words, a linear projection to $\left\langle v_{2}\right\rangle=0$ would be equivalent to a projection to $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$, where the CME backgrounds are zero. Since all those events are from a narrow centrality bin where the CME signal does not vary much, the intercept of the linear projection is presumably the CME signal. Note that there may still be complications from nonflow effects [47] depending on how $\Delta\gamma$ and $\left\langle v_{2}\right\rangle$ are measured, which is outside the scope of the present work.
@@
Since the CME background comes from the elliptic anisotropies imprinted in the POIs, it is interesting to select events with an event-shape ellipticity $v_{2}^{\rm obs}\equiv\left\langle\cos 2(\phi-\psi)\right\rangle$ calculated by POIs themselves (where the EP $\psi$ is determined in another kinematic region away from the POIs) and examine the $\Delta\gamma$ observable as a function of $v_{2}^{\rm obs}$. This method selects on the event-by-event elliptic-shape quantity (particle emissin pattern) $v_{2}^{\rm obs}$ which can fluctuate even to negative values because of large statistical fluctuations, the intercept at $v_{2}^{\rm obs}=0$, more sensitive to the CME, can be well determined from data. This event-shape (ES) analysis was performed by STAR [24] and the intercept of the linear fit is consistent with zero with the then-available statistics. However, the backgrounds are not determined by $v_{2}^{\rm obs}$ but by $\left\langle v_{\rm 2,{\rm res}}\right\rangle$. Statistically fluctuated $v_{2}^{\rm obs}=0$ does not necessarily guarantee $\left\langle v_{\rm 2,{\rm res}}\right\rangle=0$. This is illustrated by the cartoon in Fig. 1 (center). Although $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ is proportional to $\left\langle v_{2}\right\rangle$ with an unbiased event selection, the strongly (maximally) biased selection of events by $v_{2}^{\rm obs}$ will obscure the relationship between the averages $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ and the $v_{2}^{\rm obs}$ in each bin of $v_{2}^{\rm obs}$ of those events, as indicated by the hollow box in Fig. 1 (center). Because of the overall positive means $\left\langle v_{2}^{\rm obs}\right\rangle=\left\langle v_{2}\right\rangle$ and $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ over the entire event sample, the linear projection will give a positive intercept of $\left\langle v_{\rm 2,{\rm res}}\right\rangle$ at $v_{2}^{\rm obs}=0$. In other words, with finite positive average $\left\langle v_{2}\right\rangle$ in heavy ion collisions, $\left\langle v_{\rm 2,{\rm res}}\right\rangle>0$ at $v_{2}^{\rm obs}=0$ with events binned in $v_{2}^{\rm obs}$. This has been verified by model study [48]. Thus, positive residual background remains in the linear fit intercept in the analysis of $\Delta\gamma$ as a function of $v_{2}^{\rm obs}$ [24], although the background is significantly reduced compared to that in the inclusive $\Delta\gamma$ measurement. Ideally, in the ES analysis, one would want to select events with the $v_{2}^{\rm obs}$ values of all resonances (background sources) to be zero, however, such selection is experimentally insurmountable.
@@
It has been recently proposed by Xu et al. [36] to select events according to the $\hat{q}_{2,\textsc{poi}}^{2}$ of the POIs, and study the $\Delta\gamma$ as a function of $v_{2}$ of the POIs. This is similar in spirit to the ES method above where the events are selected by $v_{2}^{\rm obs}$ and the $\Delta\gamma$ is examined as a function of the same $v_{2}^{\rm obs}$ variable. Here, the selection variable $\hat{q}_{2,\textsc{poi}}^{2}$ and the $v_{2}$ variable are different, but related as they are computed by the same POIs. The $\hat{q}_{2}^{2}$ variable is an event-by-event quantity connected to two-particle cumulant anisotropy (cf Eq. 5), whereas $v_{2}$ can be measured with respect to $\psi$ from another kinematic region like in the ES method. (Of course $v_{2}$ can also be computed by the two-particle cumulant method, in which case it is more directly connected to $\hat{q}_{2}^{2}$.) The variations in the event-by-event $\hat{q}_{2,\textsc{poi}}^{2}$ quantity, within a given narrow centrality bin, are mainly from statistical fluctuations, just like the $\hat{q}_{2}^{2}$ in the ESE method. However, unlike ESE, the $v_{2}$ variable in those $\hat{q}_{2,\textsc{poi}}^{2}$-selected events by ESS are related to $\hat{q}_{2,\textsc{poi}}^{2}$ itself and are calculated from the same POIs, so the variations in $v_{2}$ is also primarily of statistical nature. (Note, we simply use $v_{2}$ here, instead of the $\left\langle v_{2}\right\rangle$ for ESE, to indicate that the $v_{2}$ in ESS is of statistical nature although it is an “average” $v_{2}$ in events selected by $\hat{q}_{2,\textsc{poi}}^{2}$.) This technique selects on the statistical fluctuations in event shape, not on the dynamical fluctuations in $\left\langle v_{2}\right\rangle$ like in the ESE method, and it is referred to as the Event-Shape-Selection (ESS) method [36]. The most important difference in the ESS method from the ESE method is that the $\hat{q}_{2}^{2}$ is computed from the POIs in the former but in the later the momentum space where to compute $\hat{q}_{2}^{2}$ is displaced from that of the POIs’. In both approaches, one groups events (within a given narrow centrality bin) according to $\hat{q}_{2}^{2}$ and studies the $\Delta\gamma$ variable as a function of the $v_{2}$ of the POIs. It is worthy to emphasize again that, while the event selection variable in both methods are dominated by statistical fluctuations event-by-event, these statistical fluctuations are canceled in $v_{2}$ in the ESE method because the calculations of $\hat{q}_{2}^{2}$ and $v_{2}$ use different sets of particles, but do not cancel (rather strongly correlated) in the $v_{2}$ in the ESS method because the calculations of $\hat{q}_{2,{\rm pair}}^{2}$ and $v_{2}$ use the same POIs.
@@ III Model setup and analysis details
Figure 2: $m_{\rm inv}$ distributions. Shown are the per-event $m_{\rm inv}$ distribution of OS pairs subtracted by that of SS pairs from avfd with axial current densities $n_{5}/s=0,0.1$, and $0.2$ (upper left), ampt (upper center), epos4 (upper right), hydjet++ (lower left), Toy Model I simulation including primordial pions, and $K_{S}$ and $\rho^{0}$-resonance decay pions (lower center), and Toy Model II simulation including in addition decay pions from a mass continuum (lower right). Particles used to form the $m_{\rm inv}$ distributions are taken from all charged hadrons (defined to be $\pi^{\pm}$, $K^{\pm}$, $p$, and $\bar{p}$ in our study and treated as pions with pion mass) and are from the kinematic range of $|\eta|<1$ and $0.2<p_{\rm T}<2$ GeV/$c$. In the toy models all final-state particles are charged pions.
@@ III.5 Toy Model I
Two versions of the toy model are examined. In Toy Model I, we include primordial pions and two resonances, $K_{S}$ and $\rho^{0}$. The resonance mass is generated according to the Breit-Wignar distribution of proper mass and width [66]. We generate particles uniformly within the rapidity range of 1.5<y<1.5-1.5<y<1.5, taking event-by-event particle multiplicities according to Poisson statistics about the means of three times the measured average pseudorapidity densities [67, 49, 68]. The $p_{\rm T}$ spectra are parameterized based on the measured data [69, 70], namely,
@@ III.6 Toy Model II
[…]